Exercise 9-14

Stellar Company uses the gross profit method to estimate inventory for monthly reporting purposes. Presented below is information for the month of May.
Inventory, May 1 $ 167,300
Purchases (gross) 580,800
Freight-in 32,100
Sales revenue 979,300
Sales returns 73,300
Purchase discounts 11,300
Compute the estimated inventory at May 31, assuming that the gross profit is 25% of net sales.
The estimated inventory at May 31 $enter the dollar amount of the estimated inventory at May 31

LINK TO TEXT

LINK TO VIDEO

Compute the estimated inventory at May 31, assuming that the gross profit is 25% of cost. (Round percentage of sales to 2 decimal places, e.g. 78.74% and final answer to 0 decimal places, e.g. 6,225.)
The estimated inventory at May 31 $enter the dollar amount of the estimated inventory at May 31
Answer #1

Beginning inventory = $167,300

Purchase = $580,800

Freight in = $32,100

Sales revenue = $979,300

Sales returns = $73,300

Purchase discount = $11,300

Gross profit = 25% of net sales

Net Sales = Sales - Sales returns

= 979,300-73,300

= $906,000

Gross profit = 25% of net sales

= 906,000 x 25%

= $226,500

Cost of goods sold = Net Sales - Gross profit

= 906,000-226,500

= $679,500

Cost of goods sold = Beginning inventory + Purchases + Freight in - Purchase discount - Ending inventory

679,500 = 167,300+580,800+32,100-11,300- Ending inventory

Ending inventory = $89,400

Gross profit = 25% of cost

Let the cost be $100

Gross profit = $25

Sales = Cost + Gross profit

= 100+25

= $125

Gross profit on sales = Gross profit/ Sales

= 25/125

= 20%

Gross profit = 20% of net sales

= 906,000 x 20%

= $181,200

Cost of goods sold = Net Sales - Gross profit

= 906,000-181,200

= $724,800

Cost of goods sold = Beginning inventory + Purchases + Freight in - Purchase discount - Ending inventory

724,800 = 167,300+580,800+32,100-11,300- Ending inventory

Ending inventory = $44,100

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks